# A Generalized Back-Door Criterion for Linear Regression

### v7

@@ Abstract
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. However, a general condition has not yet been fully established for settings in which mediators or nonlinear terms are included among the explanatory variables. To address this issue, this paper unifies Pearl’s single-door and back-door criteria and proposes a generalized criterion that enables the identification of causal effects not mediated by other explanatoly variables. In other words, for almost all parameter points in the parameter space of linear causal models, this criterion is a necessary and sufficient condition under which partial regression coefficients have a linear causal interpretation. In addition, this paper provides a quantitative characterization of post-treatment bias while controlling for confounding bias. These results are applicable to linear models with nonlinear terms, and the decomposition of linear and nonlinear causal effects is also discussed.
@@ 1 Introduction
where $X$, $M_{1}$, and $Y$ are generated by the linear structural equation model corresponding to Figure 1.
@@
As shown in (1.1), mediation analysis often makes use of post-treatment variables to block certain causal paths. When conditioning on a mediator, one might expect to control only for the mediated effect; however, it may also induce unintended bias. For example, a variable such as $M_{2}$ in Figure 1, which is influenced by multiple variables, is referred to as a collider, and it is well known that conditioning on a collider can introduce such bias. Nevertheless, the mechanism underlying post-treatment bias has not yet been fully elucidated.
@@ 2.2 Graph Terminology and Notations
An ADMG is a triple $(\bm{V},\bm{E},\bm{B})$, where $\bm{V}$ is a finite set of vertices, $\bm{E}$ is a set of directed edges that construct no directed cycles, and $\bm{B}$ is a set of bidirected edges. For vertices $a,b\in\bm{V}$, if they are connected by either a directed edge (denoted by $a\rightarrow b$) or a bidirected edge (denoted by $a\leftrightarrow b$), then $a$ and $b$ are said to be adjacent. If there is a directed edge from $a$ to $b$, then $a$ is called a parent of $b$, and $b$ is called a child of $a$. A path is a sequence of distinct vertices such that consecutive vertices are adjacent. A directed path from $a$ to $b$ (denoted by $a\rightarrow\rightarrow b$) is a path along which all edges are directed and oriented from $a$ toward $b$. A directed cycle is a concatenation of a directed path $a\rightarrow\rightarrow b$ and a directed edge $b\rightarrow a$. If there is a directed path from $a$ to $b$, then $a$ is called an ancestor of $b$, and $b$ is called a descendant of $a$. We denote the sets of ancestors and descendants of $a$ by $\bm{AN}(a)$ and $\bm{DE}(a)$, respectively. A back-door path from $a$ to $b$ (denoted by $a\dashleftarrow b$) is a path between $a$ and $b$ that begins with either a bidirected edge or a directed edge pointing into $a$. A v-structure is a path of the form $a\rightarrow c\leftarrow b$, $a\rightarrow c\leftrightarrow b$, $a\leftrightarrow c\leftarrow b$, or $a\leftrightarrow c\leftrightarrow b$, where $c$ is called a collider.
@@ 2.3 Causal Path Diagram and the Selective-Door Criterion
Let $\bm{Z}$ be a set such that $\bm{Z}\cap\{X_{i},X_{j}\}=\emptyset$. The set $\bm{Z}$ is said to block a path $P$ between $X_{i}$ and $X_{j}$ if it satisfies one of the following two conditions; otherwise, $\bm{Z}$ is said to unblock the path:
@@
    for any $X_{k}\in\bm{Z}\cap\bm{DE}(X_{j})$ such that there exists a directed path $X_{j}\rightarrow\rightarrow X_{k}$ that is unblocked by $\bm{Z}\setminus\{X_{k}\}$, every back-door path $X_{k}\dashleftarrow X_{i}$ is blocked by $(\bm{Z}\cup\{X_{j}\})\setminus\{X_{k}\}$, and
@@
In Figure 2, the set $\{M_{2}\}$ only unblocks the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\rightarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking.
@@ 2.4 Controlled Total Effects
In the literature on structural equation modeling, the decomposition of causal effects has been extensively discussed and forms the foundation of mediation analysis. In this framework, the direct effect of $X_{j}$ on $X_{i}$ is given by the path coefficient $a_{ij}$, which corresponds to the directed edge $X_{j}\to X_{i}$. The indirect effect of $X_{j}$ on $X_{i}$ is defined as the sum of products of path coefficients along each directed path $X_{j}\rightarrow\rightarrow X_{i}$, excluding the direct edge $X_{j}\to X_{i}$. The total effect of $X_{j}$ on $X_{i}$ is the sum of the direct and indirect effects. See Alwin and Hauser (1975) and Bollen (1987) for details on these causal effects.
@@
Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For a vertex set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$, consider the subgraph $\bm{G}^{\prime}$ obtained by removing all edges whose terminal vertices belong to $\bm{Z}$. Then, the $\bm{Z}$-controlled total effect of $X_{j}$ on $X_{i}$, denoted by $\tau_{ij\mid do(\bm{Z})}$, is defined as the total effect of $X_{j}$ on $X_{i}$ restricted to the subgraph $\bm{G}^{\prime}$.
@@
Assume that (ii) in Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Then, for any subset $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$, the following holds:
@@
$\tau_{ij\mid do(\bm{Z})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(X_{j}=x_{j}),do(\bm{Z}=\bm{z})],$ (2.4)
@@ 3.2 Main Results
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij\mid do(\bm{Z})}$.
@@
Corollary 3.2.
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}$ that blocks every path between $X_{j}$ and $X_{i}$. Then, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is equal to 0.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ that blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$. Then, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{j}))$ satisfying the back-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the total effect of $X_{j}$ on $X_{i}$.
@@
In many empirical analyses, it is not uncommon for all explanatory variables to be regarded as causes of interest. Thus, we also provides sufficient conditions under which all regression coefficients can be interpreted as causal effects.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on post-treatment variables (i.e., descendants in the graph) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
In addition to the set $\bm{S}=\{X_{j}\}\cup\bm{Z}$ in (3.1), the following sets are introduced:
$\bm{S}_{1}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ unblocks a back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}$
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every back-door path $X_{j}\dashleftarrow X_{i}$, the bias of the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is given by
@@
As indicated by the back-door criterion, confounding bias can be eliminated by blocking every back-door path. The first condition (i) of the selective-door criterion implies that $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. In particular, conditioning on a descendant $X_{p}$ of the outcome variable $X_{i}$ is problematic, since whenever there exists an unblocked directed path $X_{j}\rightarrow\rightarrow X_{i}$, a sequence of vertices $X_{j}\rightarrow\rightarrow X_{i}\rightarrow\rightarrow X_{p}\leftarrow\leftarrow X_{i}$ exists, leading to post-treatment bias. On the other hand, if each path from $X_{j}$ to $X_{i}$ is blocked (i.e., $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ ), the partial regression coefficient vanishes even when a decendant of the outcome variable $X_{i}$ is included among the conditioning variables.
@@ 3.4 Necessity for Almost All Parameter Points
Almost everywhere in this parameter space, satisfying the selective-door criterion is also necessary for partial regression coefficients to admit a linear causal interpretation.
@@
Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Fix a set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ and consider the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1). Then, except on a set of Lebesgue measure zero in the parameter space $\bm{\Theta}$, the following statements are equivalent:
@@ 3.5 Application to Nonlinear Data-Generating Processes
For an arbitrary error term $u_{i}$, any random vector consisting only of error terms corresponding to vertices that are not adjacent to $X_{i}$ by a bidirected edge is independent of $u_{i}$.
@@
Note that each vertices can be expressed as a function of the error terms associated with its ancestors. Depending on the nonlinear function $h$, some of the bidirected edges in Figure 3 may be omitted, since uncorrelatedness can hold. This causal path diagram shows that a bidirected edge does not necessarily indicate the presence of a confounding factor; it may also represent a causal relationship that cannot be expressed by path coefficients. For instance, the paths $X\leftrightarrow H\to Y$ and $X\to M\leftrightarrow H\to Y$ madiate the nonlinear causal effect of $X$ on $Y$.
@@
$\tau_{YX}\coloneq\frac{\partial}{\partial x}\,\mathrm{E}[Y\mid do(X=x),do(\bm{H}=\bm{h})],$

### v8

# A Unified Graphical Criterion for Characterizing a Linear Causal Interpretation of Partial Regression Coefficients

@@ Abstract
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. However, a general condition has not yet been fully established for settings in which mediators or nonlinear terms are included among the explanatory variables. To address this issue, this paper unifies d-separation, single-door, and back-door criteria and proposes a generalized graphical criterion that enables the identification of causal effects not mediated by other explanatory variables. In other words, for almost all parameter points in the parameter space of linear causal models, this criterion is a necessary and sufficient condition under which partial regression coefficients have a linear causal interpretation. In addition, this paper provides a quantitative characterization of post-treatment bias while controlling for confounding bias. These results apply to linear models represented by acyclic directed mixed graphs, where the models can include nonlinear terms. The decomposition of linear and nonlinear causal effects is also discussed.
@@ 1 Introduction
where $X$, $M_{1}$, and $Y$ are generated by the linear SEM corresponding to Figure 1.
@@
As shown in (1.1), mediation analysis often makes use of post-treatment variables to block certain causal paths. When conditioning on a mediator, one might expect to control only for the mediated effect; however, it may also induce unintended bias. For example, a variable such as $M_{2}$ in Figure 1, which is influenced by multiple variables, is referred to as a collider, and it is well known that conditioning on a collider can introduce such bias. Other structures that give rise to post-treatment bias are also discussed by Elwert and Winship (2014) and Pearl (2015). Nevertheless, the mechanism underlying post-treatment bias has not yet been fully elucidated, even when restricted to linear models.
@@ 2.2 Graph Terminology and Notations
An ADMG is a triple $(\bm{V},\bm{E},\bm{B})$, where $\bm{V}$ is a finite set of vertices, $\bm{E}$ is a set of directed edges that construct no directed cycles, and $\bm{B}$ is a set of bidirected edges. For vertices $a,b\in\bm{V}$, if they are connected by either a directed edge (denoted by $a\rightarrow b$) or a bidirected edge (denoted by $a\leftrightarrow b$), then $a$ and $b$ are said to be adjacent. If there is a directed edge from $a$ to $b$, then $a$ is called a parent of $b$, and $b$ is called a child of $a$. A path is a sequence of distinct vertices such that consecutive vertices are adjacent. A directed path from $a$ to $b$ (denoted by $a\rightarrow\rightarrow b$) is a path along which all edges are directed and oriented from $a$ toward $b$. A directed cycle is a concatenation of a directed path $a\rightarrow\rightarrow b$ and a directed edge $b\rightarrow a$. If there is a directed path from $a$ to $b$, then $a$ is called an ancestor of $b$, and $b$ is called a descendant of $a$. We denote the sets of ancestors and descendants of $a$ by $\bm{AN}(a)$ and $\bm{DE}(a)$, respectively. A back-door path from $a$ to $b$ (denoted by $a\dashleftarrow b$) is a path from $a$ to $b$ that begins with either a bidirected edge or a directed edge pointing into $a$. A v-structure is a path of the form $a\rightarrow c\leftarrow b$, $a\rightarrow c\leftrightarrow b$, $a\leftrightarrow c\leftarrow b$, or $a\leftrightarrow c\leftrightarrow b$, where $c$ is called a collider.
@@ 2.3 Causal Path Diagram and the Selective-Door Criterion
Let $\bm{Z}$ be a set such that $\bm{Z}\cap\{X_{i},X_{j}\}=\emptyset$. The set $\bm{Z}$ is said to block a path $P$ between $X_{i}$ and $X_{j}$ if it satisfies one of the following two conditions; otherwise, $\bm{Z}$ is said to open the path:
@@
    for any $X_{k}\in\bm{Z}\cap\bm{DE}(X_{j})$ such that there exists a directed path $X_{j}\rightarrow\rightarrow X_{k}$ that is opened by $\bm{Z}\setminus\{X_{k}\}$, every back-door path $X_{k}\dashleftarrow X_{i}$ is blocked by $(\bm{Z}\cup\{X_{j}\})\setminus\{X_{k}\}$, and
@@
In Figure 2, the set $\{M_{2}\}$ only opens the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\rightarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking.
@@ 2.4 Controlled Total Effects
In the literature on linear SEMs, the decomposition of causal effects has been extensively discussed and forms the foundation of mediation analysis. In this framework, the direct effect of $X_{j}$ on $X_{i}$ is given by the path coefficient $a_{ij}$, which corresponds to the directed edge $X_{j}\to X_{i}$. The indirect effect of $X_{j}$ on $X_{i}$ is defined as the sum of products of path coefficients along each directed path $X_{j}\rightarrow\rightarrow X_{i}$, excluding the direct edge $X_{j}\to X_{i}$. The total effect of $X_{j}$ on $X_{i}$ is the sum of the direct and indirect effects. See Alwin and Hauser (1975) and Bollen (1987) for details on these causal effects.
@@
Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For a vertex set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$, consider the subgraph $G^{\prime}$ obtained by removing all edges whose terminal vertices belong to $\bm{Z}$. Then, the $\bm{Z}$-controlled total effect of $X_{j}$ on $X_{i}$, denoted by $\tau_{ij\mid do(\bm{Z})}$, is defined as the total effect of $X_{j}$ on $X_{i}$ restricted to the subgraph $G^{\prime}$.
@@
Assume that (ii) in Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Then, for any subset $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$, the following holds:
@@
$\tau_{ij\mid do(\bm{Z})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}=\bm{z}),\,do(X_{j}=x_{j})],$ (2.4)
@@ 3.2 Main Results
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij\mid do(\bm{Z})}$.
@@
Corollary 3.2 (D-separation adjustment).
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every path between $X_{j}$ and $X_{i}$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is equal to 0.
@@
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ that blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
@@
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{j}))$ satisfying the back-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the total effect of $X_{j}$ on $X_{i}$.
@@
In many empirical analyses, it is not uncommon for all explanatory variables to be regarded as causes of interest. Thus, we also provide sufficient conditions under which all regression coefficients can be interpreted as causal effects.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on post-treatment variables (i.e., descendants in the causal diagram) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
In addition to the set $\bm{S}=\{X_{j}\}\cup\bm{Z}$ in (3.1), the following disjoint subsets are introduced:
$\bm{S}_{1}$ $=\{\,X_{k}\in\bm{S}\mid\text{$\bm{S}\setminus\{X_{k}\}$ opens a back-door path $X_{k}\dashleftarrow X_{i}$}.\,\}$
@@
Assume that Assumption 2.1 holds. Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ that blocks every back-door path $X_{j}\dashleftarrow X_{i}$, the bias of the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is given by
@@
As indicated by the back-door criterion, confounding bias can be eliminated by blocking every back-door path. The first condition (i) of the selective-door criterion implies that $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. In particular, conditioning on a descendant $X_{p}$ of the outcome variable $X_{i}$ is problematic, since whenever there exists an opened directed path $X_{j}\rightarrow\rightarrow X_{i}$, a sequence of vertices $X_{j}\rightarrow\rightarrow X_{i}\rightarrow\rightarrow X_{p}\leftarrow\leftarrow X_{i}$ exists, leading to post-treatment bias. On the other hand, if each path from $X_{j}$ to $X_{i}$ is blocked (i.e., $\tau_{pj\mid do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ ), the partial regression coefficient vanishes even when a decendant of the outcome variable $X_{i}$ is included among the conditioning variables.
@@ 3.4 Necessity for Almost All Parameter Points
Almost everywhere in this parameter space, satisfying the selective-door criterion is not only sufficient but also necessary for partial regression coefficients to admit a linear causal interpretation.
@@
Let $G=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Fix a set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ and consider the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1). Then, except on a set of Lebesgue measure zero in the parameter space $\bm{\Theta}$, the following statements are equivalent:
@@ 3.5 Application to Nonlinear Data-Generating Processes
Any random vector consisting only of error terms corresponding to vertices that are not adjacent to $X_{i}$ by a bidirected edge is independent of $u_{i}$.
@@
Assumption 3.8 guarantees the absence of correlation between $u_{H}$ and $u_{Y}$ when $u_{Z}$, $u_{X}$, and $u_{M}$ are independent of $u_{Y}$. Depending on the nonlinear function $h$, some of the bidirected edges in Figure 3 may be omitted, since uncorrelatedness may hold. This causal path diagram shows that a bidirected edge does not necessarily indicate the presence of a confounding factor; it may also represent a causal relationship that cannot be expressed by path coefficients. For instance, the paths $X\leftrightarrow H\to Y$ and $X\to M\leftrightarrow H\to Y$ mediate the nonlinear causal effect of $X$ on $Y$. Since model (3.3) can be described as a special case of model (2.1), as in model (3.4), all the results obtained thus far are applicable to the identification of linear causal effects.
@@
$\tau_{YX}\coloneq\frac{\partial}{\partial x}\,\mathrm{E}[Y\mid do(\bm{H}=\bm{h}),\,do(X=x)],$
@@
4 Discussion
We have derived generalized theorems that encompass existing results characterizing partial regression coefficients based on conventional graphical criteria. In causal diagrams represented by ADMGs, bidirected edges indicate the presence of unobserved common causes. In contrast, in the causal path diagrams defined in this paper, bidirected edges represent only correlations among the error terms, which constitutes a major difference. As in the case of nonlinear terms, for graphs with cycles or for partial ancestral graphs (Zhang, 2008), the results of this paper remain applicable as long as the graphs can be transformed into an ADMG while disregarding the usual interpretation of bidirected edges. This is largely due to the fact that the all proofs of the theorems in this paper rely only on algebraic arguments.
In the standard linear SEM, where bidirected edges represent the presence of unobserved confounders, the linear causal interpretation of partial regression coefficients under Theorem 3.1 can be summarized as follows:
$\beta_{ij\mid\bm{Z}}$ $=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j}),\,do(\bm{Z}^{j}=\bm{z}^{j})]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j}),\,\bm{Z}^{j}=\bm{z}^{j}]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(\bm{Z}_{j}=\bm{z}_{j}),\,do(X_{j}=x_{j})],$
where $\bm{Z}_{j}\coloneq\bm{Z}\cap\bm{DE}(X_{j})$, and $\bm{Z}^{j}\coloneq\bm{Z}\setminus\bm{Z}_{j}$. By definition, it is clear that $\tau_{ij\mid do(\bm{Z})}\equiv\tau_{ij\mid do(\bm{Z}_{j})}$. In particular, when $\bm{Z}_{j}\neq\emptyset$, special attention must be paid to the following relationship.
$\beta_{ij\mid\bm{Z}}$ $\not\equiv\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid\bm{Z}=\bm{z};\,do(X_{j}=x_{j})]$
$=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(X_{j}=x_{j}),\,\bm{Z}=\bm{z}].$
Here, $\mid\bm{Z}=\bm{z};\,do(X_{j}=x_{j})$ denotes intervening on $X_{j}$ after observing $\bm{Z}=\bm{z}$, and is conceptually similar to the counterfactual query $Q_{c}$ (Pearl, 2015). If the back-door criterion is satisfied, the symbol $\not\equiv$ can be replaced with ==. These interpretations rely on the absence of interaction or effect modification, and therefore does not necessarily hold when nonlinear terms are present. See the Appendix for details.
As a direction for future research, a more detailed investigation of the estimation of decomposed nonlinear causal effects is required. In particular, the discussion of identifiability of nonlinear causal effects in this paper is limited to whether the causal parameter (or its function) can be expressed as a function of the observed distribution. For example, the interchangeability of differentiation and expectation in (3.5) is expected to affect estimation accuracy. Therefore, further analysis under appropriate assumptions on the nonlinear terms remains an important topic for future work.